find the point estimate of the population parameter μ, and the standard deviation of the given set of data below.



Percentage of children who watch TV before bedtime.

70 67 58 60 69 62 70 62

69 59 77 59 52 79 59 59

80 42 60 59 68 40 68 68

56 66 60 40 57 57 70 71

72 54 52 67 62 59 71 72

81 49 45 78 78 69 68 69


1
Expert's answer
2021-06-03T17:19:36-0400

"\mu=\frac{\sum x_i}{n}=\frac{3039}{48}=63.3"


"\sigma=\sqrt{\frac{\sum(x_i-\mu)^2}{n}}=10.05"